# Matchings avoiding ordered patterns

### v1

@@ 3.1 Alternating paths
$e(M)\leq|[x-1]|+\frac{|[x,y]|}{2}=(x-1)+\frac{y-x+1}{2}=\frac{y+x-1}{2}<k-1$
@@
$\frac{1}{2}\left(\binom{m}{2}+4\binom{k-1}{2}\right)\leq\binom{m}{2}-\binom{m-2k+2}{2}$
@@
We learnt of simultaneous and independent work by Gaurav Kucheriya, Allan Lo, Jan Petr, Amedeo Sgueglia and Jun Yan on the Ramsey problem for the alternating path. They obtained an even stronger upper bound than Theorem 3.1, with a constant smaller than $3$ in the leading term.
@@ 3.2 Non-nested matchings
where the lower bound is conjectured to be tight [2] and the upper bound is unpublished [12]. The next lemma asserts that if the Turán question for non-nested matching behaves as we conjecture, then one obtains a better upper bound for the analogous Ramsey question.
@@ 4.1 A counter-intuitive behaviour of the interval chromatic number
The interval chromatic number $\chi_{<}(H)$ of an ordered graph $H$ is the smallest $r$ such that $V(H)$ can be partitioned into $r$ independent intervals. Pach and Tardos proved the following analogue of the Erdős–Stone–Simonovits theorem for ordered graphs.
Theorem 4.1 (Pach and Tardos).
For any ordered graph $H$ on $n$ vertices we have
$ex_{<}(n,H)=\left(1-\frac{1}{\chi_{<}(H)-1}+o(1)\right)\frac{n^{2}}{2}.$
In particular, the higher the interval chromatic number, the higher the Turán number $ex_{<}(n,H)$. The above theorem also holds if $H$ is replaced by a family of ordered graphs $\mathcal{H}$, and $\chi_{<}(H)$ is replaced by the smallest interval chromatic number among all members of $\mathcal{H}$.
Recall that we showed the following in Theorem 2.8:
@@
In light of the previous consideration on the interval chromatic number, one may think that $ex_{<}(n,\mathcal{M}^{*}_{k})$ should be (asymptotically) at most $ex_{<}(n,\text{Cross-}\mathcal{M}_{k},\text{Sep-}\mathcal{M}_{k})=(k-1)n+\Theta(k)$. Indeed, while in both cases we are forbidding a crossing $k$-matching, in the former we forbid many ordered graphs with low interval chromatic number, whereas in the latter we forbid only one additional ordered graph with arbitrary large interval chromatic number. Surprisingly, this turns out to be false: we have $ex_{<}(n,\mathcal{M}^{*}_{k})\geq\left(k-2^{-(k-2)}\right)n+\Theta(k)$.
@@
4.3 Open questions: forbidding two or three patterns
For $n\geq 2k$, we have
Another interesting question is to forbid nested, crossing and separated $k$-matchings. For $n\geq 2k$, we have
$(k-1)n\leq ex_{<}(n,\{\text{Nest-}\mathcal{M}_{k},\text{Cross-}\mathcal{M}_{k},\text{Sep-}\mathcal{M}_{k}\})\leq(k-1)n+O(k).$
The upper bound follows immediately from Theorem 2.8, whereas the lower bound follows from the following construction. Let $G$ be an ordered graph with vertex set $[n]$. Let the edge set of $G$ consists of the union of all edges incident to the vertices $\{1\}\cup[k+1,2k-2]$ and the edges lying in $2k-2]\cup\{n\}$. Then $G$ contains no separated, nested or crossing $k$-matchingand $G$ has precisely (k1)n(k-1)n edges.

### v2

@@ 3.1 Alternating paths
$e(M)\leq|[x-1]|+\frac{|[x,y]|}{2}\leq(x-1)+\frac{y-x+1}{2}=\frac{y+x-1}{2}<k-1$
@@
$\frac{1}{2}\left(\binom{m}{2}+4\binom{k-1}{2}\right)\leq 2(k-1)m-(k-1)(2k-1)$
@@
We learnt of simultaneous and independent work by Gaurav Kucheriya, Allan Lo, Jan Petr, Amedeo Sgueglia and Jun Yan on the Ramsey problem for the alternating path. They obtained an even stronger upper bound than Theorem 3.1, replacing the $3$ in the leading term with $2+\sqrt{2}/2$.
@@ 3.2 Non-nested matchings
where the lower bound is conjectured to be tight [2] and the upper bound is unpublished [12]. The next lemma asserts that if the Turán question for non-nested matching behaves as Conjecture 2.6, then one obtains a better upper bound for the analogous Ramsey question.
@@ 4 Discussion
We conclude with some final open questions and remarks. The most natural open problem is to fully settle the Turán question for non-nested $k$-matchings. Recall our main conjecture.
See 2.6
Confirming Conjecture 2.6 would immediately yield better upper bounds for the analogous Ramsey question for non-nested matchings, as showed in Lemma 3.3. In fact, it is easy to see from the proof of Lemma 3.3 that replacing the $\binom{k-1}{2}$ term in the upper bound of Theorem 3.3 with a sufficiently better constant would also yield non-trivial improvements for the Ramsey question. In the following subsections, we list some additional open problems and observations.
4.1 Turán numbers of two or three forbidden patterns
For $n\geq 2k$, we have
Question 4.1.
Is it the case that $ex_{<}(n,\{\text{Nest-}\mathcal{M}_{k},\text{Cross-}\mathcal{M}_{k}\})=ex_{<}(n,\neg\text{Sep-}\mathcal{M}_{k})$?
Another interesting question is to determine the maximum number of edges when forbidding a nested, crossing and separated $k$-matchings. For $n\geq 2k$ and $k\geq 3$, we have
$(k-1)n\leq ex_{<}(n,\{\text{Nest-}\mathcal{M}_{k},\text{Cross-}\mathcal{M}_{k},\text{Sep-}\mathcal{M}_{k}\})\leq(k-1)n+O(k^{3}).$
The upper bound follows immediately from Theorem 2.9, whereas the lower bound follows from the following construction. Let $G$ be an ordered graph with vertex set $[n]$. Let the edge set of $G$ consists of the union of all edges incident to the vertices $\{1\}\cup[k+1,2k-2]$ and the edges lying in $2k-2]\cup\{n\}$. Then $G$ contains no separated, nested or crossing $k$-matching (for $k\geq 3$) and $G$ has precisely (k1)n(k-1)n edges. This could be tight.
Question 4.2.
Do we have $ex_{<}(n,\{\text{Nest-}\mathcal{M}_{k},\text{Cross-}\mathcal{M}_{k},\text{Sep-}\mathcal{M}_{k}\})=(k-1)n$ for $k\geq 3$?
@@ 4.2 A counter-intuitive behaviour of the interval chromatic number
Theorem 1.2 by Pach and Tardos states that the interval chromatic number $\chi_{<}(H)$ governs the asymptotic behaviour of $ex_{<}(n,H)$. In particular, if $\chi_{<}(H_{1})<\chi_{<}(H_{2})$ then (asymptotically) we have $ex_{<}(n,H_{1})>ex_{<}(n,H_{2})$. Recall that we showed the following in Theorem 2.9:
@@
In light of the previous consideration on the interval chromatic number, one may think that $ex_{<}(n,\mathcal{M}^{*}_{k})$ should be (asymptotically) at most $ex_{<}(n,\text{Cross-}\mathcal{M}_{k},\text{Sep-}\mathcal{M}_{k})=(k-1)n+\Theta(k)$. Indeed, while in both cases we are forbidding a crossing $k$-matching, in the former we forbid many ordered graphs with low interval chromatic number, whereas in the latter we forbid only one additional ordered graph with arbitrary large interval chromatic number. Perhaps surprisingly, this turns out to be false: we have $ex_{<}(n,\mathcal{M}^{*}_{k})\geq\left(k-2^{-(k-2)}\right)n+\Theta(k)$.